OpenStudy (anonymous):

five-sixths of a number is -60 what is the number?

OpenStudy (amistre64):

5/6 is just another form of percent

myininaya (myininaya):

\[\frac{5}{6}n=-60\]

myininaya (myininaya):

solve for n! :)

OpenStudy (amistre64):

N(5/6) = -60 N = -60(6/5)

myininaya (myininaya):

by the way i don't mean that as factorial

OpenStudy (amistre64):

n*n-1*n-2*n-3*...*3*2 lol

myininaya (myininaya):

lol

myininaya (myininaya):

of is multiply is is equal